What mass of H2O is required to react with 1.5 g of Na2S2O3 ?
Step 2: Mass to moles
Step 3: Mole to Mole (Ratio)
Step 4: Mole to Mass
1
Expert's answer
2020-12-06T13:57:33-0500

The balanced reaction equation of cloth bleach using Na2S2O3 and chlorine is:

Na2S2O3(aq) + 4Cl2(g) + 5H2O(aq) "\rightarrow" 2NaHSO4(aq) + 8HCl(aq)

The number of the moles of Na2S2O3 is its mass (1.5 g) divided by its molar mass:

158.11 g/mol:

"n = \frac{m}{M} = \frac{1.5}{158.11} = 0.0095" mol.


According to the reaction equation, one molecule of Na2S2O3 reacts with five molecules of H2O:

"n(Na_2S_2O_3) = \frac{n(H_2O)}{5}"

"n(H_2O) = 5\u00b7n(Na_2S_2O_3) = 5\u00b70.0095 = 0.047" mol.

Therefore, the mass of water needed is:

"m(H_2O)= nM = 0.047\u00b718.02 = 0.855" g.


Answer: 0.855 g of H2O is required to react with 1.5 g of Na2S2O3.